Circular No. 6374

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) H. A. Weaver, Applied Research Corporation; P. D. Feldman, Johns

Hopkins University; M. F. A’Hearn, University of Maryland; C. Arpigny, University of Liège; and J. C. Brandt and C. E. Randall, University of Colorado, report on their Hubble Space Telescope spectroscopic observations: “Ultraviolet spectra (range 129.5–327.7 nm) were obtained on Apr.

1. The region below 220.0 nm is dominated by emission from CO. At least 25 bands of the CO Fourth Positive Group were detected, including bands in the progression v′ = 14 pumped via the solar H Lyα line. Seven or more bands of the forbidden CO Cameron band system were detected. These latter bands are produced primarily by prompt emission following the photodissociation of CO_{2} and provide a tracer of CO_{2} in the nucleus. Preliminary calculations indicate that the molecular production rates for both CO and CO_{2} were ∼ 1–2 × 10^{28} mol/s, which corresponds to an abundance relative to water of ∼ 5 percent. The S_{2} molecule was clearly detected via fluorescence in multiple bands between 285.0 and 312.0 nm, at a relative abundance of roughly 0.01 percent. This is the first definite detection of S_{2} in a comet since its initial discovery during IUE observations of C/1983 H1 (IRAS-Araki-Alcock) in 1983 (cf. IAUC 3809). The C/1996 B2 spectra have several as yet unidentified emissions, which we are currently investigating.”

M. A. DiSanti, M. J. Mumma, N. Dello Russo, and D. X. Xie, Goddard Space Flight Center, NASA; M. Fomenkova, University of California, San Diego; and K. Magee-Sauer, Rowan College of New Jersey, report observations of various parent molecules in C/1996 B2 on Mar. 24 UT, using the CSHELL cryogenic infrared spectrometer at the NASA Infrared Telescope Facility, Mauna Kea: “From the line flux contained within a 1′′ × 1′′.4 aperture centered on the nucleus, the production rates (in molecules/s) are: Q(H_{2}O) = 1.7 × 10^{29}, revised slightly from IAUC 6366; Q(CO) = 9.8 × 10^{27}, from analysis of the P2, P3 lines of the 1-0 band; Q(CH_{4}) = 1.2 × 10^{27}, from the R0 line of the ν_{3} band; and Q(C_{2}H_{6}) = 6.4 × 10^{26}, based on the ^{r}Q_{0} and ^{r}Q_{1} branches of the ν_{7} band. The relative production rates for these parent volatiles are thus 100:5.8:0.7:0.4. The CH_{4}-production rate assumes a rotational temperature of 50 K; for T = 100 K, Q(CH_{4}) would be ∼ 2.5 times larger. The Q(C_{2}H_{6}) value represents only the population in those K-ladders with K = 0-3; thus the production rate could be up to ∼ 3 times larger, if C_{2}H_{6} is characterized by a temperature as high as 130 K. Our production rates are sensitive to the direct nuclear source, primarily.”

1996 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green